SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Friday, May 27, 2005

Status Report From: China National Space Administration
Posted: Saturday, May 28, 2005

  DAY 147

DMSP

OPS-49 Rev 28175 R+05 at POGO-C at 10:41z on May 26: Unable to get 2k at rise with POGO-C. POGO issued track order. CAPS still not receiving telemetry. CAPS configured EPATH1 and still did not receive 2K. Directed POGO to configure for ACS support and CAPS reconfigured to connect through the AIU. CAPS received 2K through the AIU and recovered all but 35 sec of the smooth data. 2 fine data sets were pushed back to the next rev and recovered in full. USER declared Cat III data loss of 1437 sc seconds /35.92 pbk seconds.

All DMSP operations were nominal over the past 24 hours.

POES

*NOAA-18 Update

SPACECRAFT ACTIVITIES

  • Placed MHS in scan mode
  • Slewed the solar array to negative 35 degrees, (-35)

ACTIVITIES SCHEDULED FOR NEXT 24 HOURS

  • Turn on DCS
  • Turn on SBUV

All POES operations were nominal over the past 24 hours.

GOES

All other GOES operations were nominal over the past 24 hours.

Questions or comments to Keith.Amburgey@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.